# Stationary densities and delocalized domain walls in asymmetric exclusion processes competing for finite pools of resources

### v1

@@ VII CONCLUSION
In this work, we have investigated the stationary densities and domain walls in a model composed of two TASEP lanes connected antiparallelly to two particles reservoirs. This was originally proposed in Ref. astik-erwin as a model for TASEPs with distributed resources. Considering the same model but in a different reduced parameter space, where the entry rates in both the TASEP lanes are equal and parametrized by $\alpha$ and the exit rates are in both lanes are also equal and parametrized by $\beta$. In addition, two functions $f$ and $g$ are defined, which control the effective entry and exit rates, respectively. The exact quantitative results depend upon the forms of the functions $f$ and $g$ selected here, but even if $f$ and $g$ have different functional forms, the results will still be qualitatively similar as long as they are increasing and decreasing monotonically with their arguments and the resources are finite. To keep things simple, the functions $f$ and $g$ are connected in a straightforward way and are presumed to have simple forms. This guarantees that $f$ increases monotonically with the reservoir population while $g$ decreases monotonically. We study the models by employing mean-field theories, supplemented by extensive Monte Carlo simulations. We calculate the stationary densities and the phase diagrams in terms of $\alpha,\beta$ and filling factor $\mu$.
@@
In spite of the simplifying reduction of the model, it shows hitherto unknown and complex phase behavior. For instance, the two TASEP lanes have have the same phases — LD-LD, HD-HD, MC-MC and DW-DW; distinct phases in the two TASEP lanes are not possible, unlike in the work of Ref. astik-erwin . This forms a major conclusion from this work. Furthermore, the phase diagrams are very different from those reported in Ref. astik-erwin . Remarkably, in the DW-DW phase this model admits domain walls in each of the TASEP lane only in the form of a pair of delocalized domain walls (DDWs), and not a single localized domain wall (LDW). Rather interestingly, the phase space region occupied by the DDW solutions covers an extended region. This is in stark contrast to other known models of TASEP, whether with open boundaries, or closed geometries or finite resources such as the one studied in Ref. astik-erwin . Indeed, all such other models of TASEP, the phase space region for a single or more than one DDWs is inevitably a line in the plane of the control parameters. This is not just a theoretical novelty, it is strong phenomenological significance as well. For instance, in possible practical realizations of this model, the presence of a pair of DDWs in an extended region of the phase space means that for an extended range of the control parameters (here, $\alpha,\beta$), the stationary densities are actually strongly fluctuating even in the thermodynamic limit (see Figs. 4 and 5), since the DDWs necessarily involve mass fluctuations that scale with the system size (and hence survive in the thermodynamic limit). Considering this as a simplified description for ribosome translation along messenger RNA strands in biological cells with a fixed number of ribosomes available, we expect for some ranges of the entry and exit rate parameters, there should be large fluctuations of the ribosome numbers along the messenger RNA strands. We have further studied the phase transitions, all of which are continuous in nature, with the bulk density appearing as the order parameter. We have also calculated all of the phase boundaries analytically within mean-field theory, which are in good agreement with the corresponding MCS results. We have also identified the different threshold values of $\mu$ at which different phase transitions occur in this model (see Table 2). We have also calculated the occupations $N_{1}$ and $N_{2}$ of the two reservoirs, which are always equal, independent of the phases on the TASEP lanes. Thus the model in this particular reduced parameter space chosen here does not allow any population imbalances in different parts of a bulk system.
@@
We now briefly compare and contrast our results with those in Ref. astik-erwin . While the two models are structurally same and both admit the particle-hole symmetry, the reduced parameter space chosen here is distinct from the one considered in Ref. astik-erwin . As a result, the phase diagrams in the present study are quite different from those reported in Ref. astik-erwin . The two TASEP lanes have statistically identical stationary densities here, unlike in Ref. astik-erwin , where distinct stationary densities on the two TASEP lanes are also possible. Related to this, the two reservoirs always have the same steady state occupations in the present study, ruling out any macroscopic population imbalances, which differs from Ref. astik-erwin . Furthermore, all the phase boundaries in this study correspond to continuous transitions, whereas discontinuous transitions were also found in Ref. astik-erwin . Lastly and as mentioned above, the present model admits a pair of DDWs over extended regions of the parameter space, whereas for model in Ref. astik-erwin , a pair of DDWs are possible only for the control parameters falling on lines in the phase space. On the other hand, there are no localized domain walls in the present model, which are found in Ref. astik-erwin for the control parameters in extended regions of the phase space.
@@
This work can be extended in various ways. For instance, one can increase the number of TASEP lanes and number of reservoirs. Our above framework including our mean-field analysis can be easily extended to any number of reservoirs and TASEP lanes with equal entry and exit rates (as here). Nonetheless, we expect the broad qualitative features observed in our study to persist. For instance, it is reasonable to expect that all the channels are to be in their LD (HD) phases for very low (high) resources $\mu$. As $\mu$ rises DW and MC phases should appear. It is expected that these DWs should all be delocalized for reasons similar to the above. One can also consider using reservoirs having functions defining the entry and exit rates that are different from the ones used here. These qualitative predictions can be confirmed by stochastic simulation studies. Furthermore, making the parameters $a,b$, which parametrize the effective entry and exit rates, fluctuate stochastically with the specified distributions would be still another intriguing extension. Our study here should serve as a catalyst for further investigation in these directions.
@@ Appendix A Particle-hole symmetry and phase diagrams for $1<\mu\leq 2$
We now demonstrate that the particle-hole symmetry is reflected in the phase diagrams of our model. In the main text, we presented phase diagrams for $\mu\leq 1$; see Fig. 2. Here, we extend the analysis to the complementary regime $1<\mu\leq 2$, which—by symmetry—can be inferred from the previously obtained phase diagrams corresponding to $\mu\leq 1$ via the transformations (94)–(96). Figures 8(a) ($\mu=9/8$), 8(b) ($\mu=11/8$), and 8(c) ($\mu=13/8$) are related, respectively, to Figs. 2(c) ($\mu=7/8$), 2(b) ($\mu=5/8$), and 2(a) ($\mu=3/8$) by the transformations (94)–(96).
@@ Appendix B Steady state density profiles in the LD-LD, HD-HD, and MC-MC phases
Stationary state density profiles in the LD-LD, HD-HD, and MC-MC phases are presented in Fig. 9.

### v2

@@ VII CONCLUSION
In this work, we have investigated the stationary densities and domain walls in a model composed of two TASEP lanes connected antiparallelly to two particle reservoirs. This was originally proposed in Ref. [53] as a model for TASEPs with distributed resources. Considering the same model but in a different reduced parameter space, where the entry rates in both the TASEP lanes are equal and parametrized by $\alpha$ and the exit rates are in both lanes are also equal and parametrized by $\beta$. In addition, two functions $f$ and $g$ are defined in Eq. (4), which control the effective entry and exit rates, respectively. The exact quantitative results depend upon the forms of the functions $f$ and $g$ selected here, but even if $f$ and $g$ have different functional forms, the results will still be qualitatively similar as long as they are increasing and decreasing monotonically with their arguments and the resources are finite. To keep things simple, the functions $f$ and $g$ are connected in a straightforward way and are presumed to have simple forms. This guarantees that $f$ increases monotonically with the reservoir population while $g$ decreases monotonically. We study the models by employing mean-field theories, supplemented by extensive Monte Carlo simulations. We calculate the stationary densities and the phase diagrams in terms of $\alpha,\,\beta$ and filling factor $\mu$.
@@
In spite of the simplifying reduction of the model, it shows hitherto unknown and complex phase behavior. For instance, the two TASEP lanes have the same phases — LD-LD, HD-HD, MC-MC and DW-DW; distinct phases in the two TASEP lanes are not possible, unlike in the work of Ref. [53]. This forms a major conclusion from this work. Furthermore, the phase diagrams are very different from those reported in Ref. [53]. Remarkably, in the DW-DW phase this model admits domain walls in each of the TASEP lane only in the form of a pair of delocalized domain walls (DDWs), and not a single localized domain wall (LDW). Rather interestingly, the phase space region occupied by the DDW solutions covers an extended region. This is in stark contrast to other known models of TASEP, whether with open boundaries, or closed geometries or finite resources such as the one studied in Ref. [53]. Indeed, all such other models of TASEP, the phase space region for a single or more than one DDWs is inevitably a line in the plane of the control parameters. Interestingly, the large number fluctuations in the TASEP lanes in the DW-DW phase is not accompanied by a corresponding large fluctuations in the reservoir particles numbers; the relative fluctuations in the latter actually should vanish in the thermodynamic limit. This is not just a theoretical novelty, it has a strong phenomenological significance as well. For instance, in possible practical realizations of this model, the presence of a pair of DDWs in an extended region of the phase space means that for an extended range of the control parameters (here, $\alpha,\,\beta$), the stationary densities are actually strongly fluctuating even in the thermodynamic limit (see Figs. 6 and 7), since the DDWs necessarily involve mass fluctuations that scale with the system size (and hence survive in the thermodynamic limit). Considering this as a simplified description for ribosome translation along messenger RNA strands in biological cells with a fixed number of ribosomes available, we expect for some ranges of the entry and exit rate parameters, there should be large fluctuations of the ribosome numbers along the messenger RNA strands. While any direct experimental results on such large fluctuations are not yet known, we expect our studies will provide new impetus to study along these lines. We have further studied the phase transitions, all of which are continuous in nature, with the bulk density appearing as the order parameter. We have also calculated all of the phase boundaries analytically within mean-field theory, which are in good agreement with the corresponding MCS results. Moreover, we have also identified the different threshold values of $\mu$ at which different phase transitions occur in this model (see Table 2). Furthermore, we have calculated the occupations $N_{1}$ and $N_{2}$ of the two reservoirs, which are always equal, independent of the phases on the TASEP lanes. Thus the model in this particular reduced parameter space chosen here does not allow any population imbalances in different parts of a bulk system.
@@
We now briefly compare and contrast our results with those in Ref. [53]. While the two models are structurally same and both admit the particle-hole symmetry, the reduced parameter space chosen here is distinct from the one considered in Ref. [53]. As a result, the phase diagrams in the present study are quite different from those reported in Ref. [53]. The two TASEP lanes have statistically identical stationary densities here, unlike in Ref. [53], where distinct stationary densities on the two TASEP lanes are also possible. Related to this, the two reservoirs always have the same steady state occupations in the present study, ruling out any macroscopic population imbalances, which differs from Ref. [53]. Furthermore, all the phase boundaries in this study correspond to continuous transitions, whereas discontinuous transitions were also found in Ref. [53]. Lastly and as mentioned above, the present model admits a pair of DDWs over extended regions of the parameter space, whereas for model in Ref. [53], a pair of DDWs are possible only for the control parameters falling on lines in the phase space. On the other hand, there are no localized domain walls in the present model, which are found in Ref. [53] for the control parameters in extended regions of the phase space.
We note that in spite of the adhoc nature of the MFT, it agrees very well with the MCS results, a fact also holds for Ref. [53]. This is not generally expected as the overall particle number conservation should make the density fluctuations correlated, which is neglected in MFT. This unexpected agreement between the MFT and MCS results may be heuristically understood by considering that the presence of the attached particle reservoirs ensure that there are no conservation laws associated with the individual TASEP lanes, making them similar to open TASEPs. The latter have no conservation laws, and consequently, the density fluctuations are uncorrelated, making the MFT a good and accurate described. The presence of the particle reservoirs similarly make the density fluctuations in the TASEP lane uncorrelated, which makes the MFT accurate. See also Ref. [26] similar discussions in a related context.
@@
This work can be extended in various ways. For instance, one can increase the number of TASEP lanes and number of reservoirs. Our above framework including our mean-field analysis can be easily extended to any number of reservoirs and TASEP lanes with equal entry and exit rates (as here). Nonetheless, we expect the broad qualitative features observed in our study to persist. For instance, it is reasonable to expect that all the channels are to be in their LD (HD) phases for very low (high) availability of resources $\mu$. As $\mu$ rises DW and MC phases should appear. It is expected that these DWs should all be delocalized for reasons similar to the above. One can also consider using reservoirs having functions defining the entry and exit rates that are different from the ones used here. These qualitative predictions can be confirmed by stochastic simulation studies. Furthermore, making the parameters $\alpha,\,\beta$, which parametrize the effective entry and exit rates, fluctuate stochastically with the specified distributions would be still another intriguing extension. Our study here should serve as a catalyst for further investigation in these directions.
@@ Appendix A Particle-hole symmetry and phase diagrams for $1<\mu\leq 2$
We now demonstrate that the particle-hole symmetry is reflected in the phase diagrams of our model. In the main text, we presented phase diagrams for $\mu\leq 1$; see Fig. 2. Here, we extend the analysis to the complementary regime $1<\mu\leq 2$, which—by symmetry—can be inferred from the previously obtained phase diagrams corresponding to $\mu\leq 1$ via the transformations (83)–(85). Figures 10(a) ($\mu=9/8$), 10(b) ($\mu=11/8$), and 10(c) ($\mu=13/8$) are related, respectively, to Figs. 2(c) ($\mu=7/8$), 2(b) ($\mu=5/8$), and 2(a) ($\mu=3/8$) by the transformations (83)–(85).
@@ Appendix B Steady state density profiles in the LD-LD, HD-HD, and MC-MC phases
Stationary state density profiles in the LD-LD, HD-HD, and MC-MC phases are presented in Fig. 11. Both MFT and MCS studies results are shown, which mutually agree well.
@@
Appendix C Detailed calculations of the HD-HD phase
In this section, we explicitly derive the effective entry and exit rates in the HD-HD phase, determine the extent of the HD-HD region in the $\alpha$-$\beta$ plane, and obtain the stationary density characterizing the HD-HD phase.
The HD-HD phase also corresponds to the solution $\rho^{(1)}=1-\beta_{\text{eff}}^{(1)}=\rho^{(2)}=1-\beta_{\text{eff}}^{(2)}\equiv\rho_{\text{HD-HD}}>1/2$. This yields $N_{1}=N_{2}$ similar to the LD-LD phase:
$1-\beta_{\text{eff}}^{(1)}=1-\beta_{\text{eff}}^{(2)}$
$\implies\quad$ $N_{1}=N_{2}.$ (87)
$\frac{N_{1}}{L}=\frac{\mu-1+\beta}{1+\beta}=\frac{N_{2}}{L}.$ (88)
Again, $N_{1}$ and $N_{2}$ scale linearly with the system size $L$, and depend linearly on the filling factor $\mu$.
Since $0<N_{1},N_{2}\leq L$, Eq. (88) yields
$\mu\leq 2\quad\text{and}\quad\frac{2-\mu}{1+\beta}<1.$ (89)
$\frac{2-\mu}{1+\beta}<\frac{\alpha}{\alpha+\beta},$ (92)
$\frac{(2-\mu)\beta}{1+\beta}<\frac{1}{2}.$ (93)
Inequalities (89), (92), and (93) together define the HD-HD region in the $\alpha$-$\beta$ plane for a fixed $\mu$. These conditions are identical to those obtained via particle-hole symmetry, Eqs. (30), (31), and (32).
$\rho_{\text{HD-HD}}=1-\beta_{\text{eff}}^{(1(,2))}=\frac{1-\beta+\mu\beta}{1+\beta}.$ (94)